OpenStudy (anonymous):

The hypotenuse of a right triangle is 50 mm long. One leg of the right triangle is 30 mm long. What is the length of the other leg? twenty millimeters thirty millimeters forty millimeters sixty millimeters

OpenStudy (triciaal):

this is a 3,4 5 triangle using pythagorean theorem and letting the unknown leg = x 50^2 - 30^2 = x^2 (50-30)(50+ 30) = x^2 20 *80 = x^2 1600 = x^2 x = 40

OpenStudy (triciaal):